# DM-MPPI: Datamodel for Efficient and Safe Model Path Integral Control

### v1

@@ III-C Online Inference Phase
Given a new set of $K$ sampled trajectories, we first compute the cost components $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ and total costs $C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$. We then compute the instance statistics
$\bar{C}=\frac{1}{K}\sum_{k=1}^{K}C_{k},\qquad\sigma_{C}=\text{std}(\{C_{k}\}_{k=1}^{K})$ (19)
and obtain the predicted influence coefficients as
$\hat{\theta}_{k}=h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ (20)
@@
Sample Pruning for Efficiency. We retain only those samples whose predicted influence magnitude exceeds a threshold $\tau$:
$\mathcal{K}_{\text{keep}}=\{k:|\hat{\theta}_{k}|\geq\tau\}$ (21)
This pruning step reduces computational cost while preserving the samples that matter most. The MPPI control update is then computed using only the retained samples:
@@
Adaptive Constraint Penalty for Safety. To adapt the constraint penalty online, we monitor the influence contributed by constraint-violating samples. We define the violation influence ratio as
$r_{\text{viol}}=\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$ (23)
where $\mathcal{K}_{\text{viol}}=\{k:C_{\text{viol},k}>0\}$ denotes the set of constraint-violating samples. This ratio quantifies how much of the total influence comes from samples that violate constraints.
When $r_{\text{viol}}$ is high, constraint-violating trajectories disproportionately affect the control output, suggesting the current penalty $\rho$ is insufficient to discourage unsafe behaviors. We update the penalty coefficient according to
$\rho_{t+1}=\rho_{t}+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$ (24)
where $\eta>0$ is a step size and $r_{\text{target}}\in(0,1)$ is the desired maximum violation influence ratio. This adaptation increases $\rho$ when violation influence exceeds the target and decreases it otherwise, maintaining constraint satisfaction while avoiding overly conservative behavior. Algorithm 2 presents the complete online procedure.
@@
0: Trained predictor $h_{\phi^{*}}$, initial state $x_{0}$, nominal control $\bar{U}$, initial penalty $\rho$, pruning threshold $\tau$, target violation ratio $r_{\text{target}}$, adaptation step size $\eta$, temperature $\lambda$
1:for each control iteration do
2:  Sample $K$ trajectories from current state
3:  Compute cost components: $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ for all $k$
4:  Compute total costs: $C_{k}\leftarrow C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$
5:  Compute instance statistics: $\bar{C}\leftarrow\frac{1}{K}\sum_{k}C_{k}$,  $\sigma_{C}\leftarrow\text{std}(\{C_{k}\})$
6:  Predict influence: $\hat{\theta}_{k}\leftarrow h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ for all $k$
7:  Identify kept samples: $\mathcal{K}_{\text{keep}}\leftarrow\{k:|\hat{\theta}_{k}|\geq\tau\}$
8:  Identify violating samples: $\mathcal{K}_{\text{viol}}\leftarrow\{k:C_{\text{viol},k}>0\}$
9:  Compute violation ratio: $r_{\text{viol}}\leftarrow\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$
10:  Update penalty: $\rho\leftarrow\rho+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$
11:  Compute weights: $\tilde{w}_{k}\leftarrow\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ for $k\in\mathcal{K}_{\text{keep}}$
12:  Compute optimal control: $u^{*}\leftarrow\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k}$
@@
Remark 1 (Unified Influence for Efficiency and Safety).
Because the predictor takes both total cost and violation cost as input, a single learned model captures influence relevant to both sample efficiency and constraint handling. The pruning mechanism improves efficiency by discarding low-influence samples regardless of their constraint status, while the adaptive $\rho$ mechanism improves safety by strengthening penalties when constraint-violating samples remain influential. These two mechanisms operate complementarily: pruning reduces computation without sacrificing control quality, while penalty adaptation ensures the kept samples lead to safe behaviors.
Remark 2 (Computational Complexity).
The offline phase requires $O(N\cdot M\cdot K)$ subset evaluations and $N$ LASSO solves, but is performed only once. The online phase requires only $O(K)$ forward passes through $h_{\phi^{*}}$ per control iteration, making influence prediction negligible compared to trajectory sampling and cost evaluation. The pruning step further reduces the cost of the weighted average computation from $O(K)$ to $O(|\mathcal{K}_{\text{keep}}|)$.
@@ IV Experiments
We evaluate the proposed framework on a path tracking task with obstacle avoidance, demonstrating unified benefits for sample efficiency and constraint-aware control.
IV-A Setup
We consider a vehicle governed by the kinematic bicycle model with state $x=[p_{x},p_{y},\psi,v]^{\top}$ (position, heading, velocity) and control $u=[\delta,a]^{\top}$ (steering angle, acceleration). The vehicle tracks an oval reference path while avoiding circular obstacles. The goal cost penalizes deviation from the reference path, while the violation cost applies a quadratic penalty when the vehicle enters the obstacle regions.
The MPPI controller uses a planning horizon of $T=20$ steps at $dt=0.1$ s and temperature $\lambda=100$. Control noise is sampled from $\mathcal{N}(0,\Sigma)$ with $\Sigma=\text{diag}(0.1,0.5)$ for steering and acceleration respectively.
For the influence predictor $h_{\phi}$, we use a multilayer perceptron (MLP) with two hidden layers of 64 units each and ReLU activations. The network takes as input the normalized costs $(S_{k},S_{\text{viol},k},\bar{S},\sigma_{S})$ and outputs the predicted influence coefficient $\hat{\theta}_{k}$. Training uses the Adam optimizer with learning rate $10^{-3}$ for 1000 epochs. The datamodel parameters are set as: inclusion probability $\alpha=0.5$, number of subsets $M=50$, and LASSO regularization $\mu=0.01$. We collect $N=200$ MPPI instances from randomly sampled initial states for offline training.
IV-B Sample Efficiency Analysis
Figure 2: Sample efficiency comparison between DM-MPPI and standard MPPI across varying sample sizes $K\in\{50,100,150,\ldots,500\}$.
To determine the minimum number of samples required for effective control, we evaluate both methods across sample sizes $K\in\{50,100,150,\ldots,500\}$. Figure 2 presents the results. Standard MPPI shows a steep performance degradation below $K=200$, with position RMSE increasing from 0.41 m at $K=500$ to 1.30 m at $K=50$. In contrast, DM-MPPI maintains stable performance across a wider range: at $K=100$, it achieves position RMSE of 0.44 m and $r_{\text{viol}}=0.11$, comparable to standard MPPI at $K=500$ (0.41 m and 0.09 respectively).
However, the datamodel approach has a minimum sample requirement. At $K=50$, DM-MPPI performance degrades noticeably (position RMSE increases to 0.72 m, $r_{\text{viol}}$ rises to 0.35), suggesting insufficient samples for reliable influence estimation. This aligns with the effective influence size $K_{\text{eff}}^{\theta}\approx 28$ observed in our analysis—when the total sample count approaches this threshold, the datamodel cannot effectively distinguish high-influence samples from noise.
Based on this analysis, we select $K=100$ for DM-MPPI in subsequent experiments, achieving $5\times$ sample reduction compared to standard MPPI ($K=500$) while maintaining equivalent control quality.
IV-C Performance Comparison
Figure 3: Comparison of sampled different trajectories.
TABLE I: Performance Comparison
Metric Standard DM-Fixed DM-Adaptive
Samples used 500 100 100
Position RMSE (m) 0.41 0.44 0.43
Heading RMSE (rad) 0.031 0.033 0.032
Min. obstacle dist. (m) 1.24 1.31 1.42
$r_{\text{viol}}$ 0.091 0.082 0.071
Iteration time (ms) 85.4 18.2 18.5
We compare three configurations: (1) Standard MPPI with $K=500$ samples and fixed penalty $\rho=10^{10}$, (2) DM-MPPI with $K=100$ samples and fixed $\rho=10^{10}$, and (3) DM-MPPI with $K=100$ samples and adaptive $\rho$ using $r_{\text{target}}=0.05$ and step size $\eta=10^{9}$.
Table I summarizes the performance comparison. The datamodel approach enables $5\times$ sample reduction while maintaining tracking quality. Figure 3 visualizes the sampled trajectories at a representative timestep; despite using only 100 samples, the datamodel variant produces a control output comparable to standard MPPI with 500 samples. Analysis of the learned influence coefficients reveals an effective influence size of $K_{\text{eff}}^{\theta}\approx 28$, confirming that roughly one-quarter of the samples carry meaningful influence while the remainder can be safely pruned. The adaptive variant achieves the largest minimum obstacle distance by dynamically adjusting $\rho$ based on the violation influence ratio $r_{\text{viol}}$.
V Limitations and Future Work
Soft Constraint Guarantees. The proposed framework relies on soft constraint penalties, which cannot guarantee hard constraint satisfaction. While the adaptive $\rho$ mechanism reduces the influence of constraint-violating samples, it cannot ensure that the resulting control will never violate constraints. Safety-critical applications may require integration with formal methods such as Control Barrier Functions [20, 21] or tube-based robust MPC [17]. Future work could integrate the datamodel framework with CBFs, using predicted influence to inform barrier parameters.
Limited Experimental Scope. The current evaluation considers a single path tracking task with obstacle avoidance using a kinematic bicycle model. While this serves as a representative benchmark, comprehensive validation requires evaluation across diverse domains including manipulation, legged locomotion, and higher-dimensional systems with more complex constraint structures. Future work will include comparisons against state-of-the-art sample-efficient MPPI variants [7, 16], evaluation under varying noise levels and constraint complexities, and deployment on real hardware platforms.
Fixed Predictor After Offline Training. The influence predictor is fixed after offline training, which limits adaptability to changing conditions. Online fine-tuning would allow the predictor to adapt to distribution shifts or novel scenarios, but introduces computational overhead that conflicts with real-time requirements. One promising direction is to maintain a small replay buffer of recent MPPI instances and perform incremental updates during idle cycles. Alternatively, meta-learning approaches could train the predictor to rapidly adapt to new conditions with minimal fine-tuning samples.
Generalization Across Cost Scales. The framework uses a fixed inclusion probability $\alpha$ during offline training, and the learned predictor generalizes well within the range of cost distributions encountered during training. However, performance may degrade when deployed in scenarios where cost statistics differ substantially from training. Normalizing costs by instance-level statistics partially addresses this issue, but does not fully resolve it when the underlying cost structure changes significantly. Future work could explore scale-invariant predictor architectures that operate on relative cost rankings rather than absolute values to enable broader generalization.
VI Conclusion
We presented a framework that extends the datamodeling paradigm from machine learning to sampling-based optimal control. While traditional datamodels estimate sample influence through regression on a fixed dataset, our approach trains an influence predictor that generalizes across control instances—enabling real-time influence estimation without solving regression problems online. By learning to map cost features to influence coefficients, the predictor identifies which samples contribute meaningfully to the control output, allowing efficient sample pruning and adaptive constraint handling through a single learned model.
The key insight is that influence coefficients, though computed via subset regression during offline training, show predictable structure as a function of sample costs and global statistics. This structure enables amortized inference: rather than re-running regression for each new MPPI instance, we predict influence directly from cost features in a single forward pass. Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining tracking performance and achieving improved safety margins through adaptive penalty tuning.

### v2

@@ III-C Online Inference Phase
Given a new set of $K$ sampled trajectories, we compute cost components and build the 8-dimensional feature vector $\mathbf{f}_{k}$ for each sample as in (13). The predicted influence magnitudes are
$\hat{\theta}_{k}=h_{\phi^{*}}(\mathbf{f}_{k})$ (15)
@@
Sample Pruning for Efficiency. We retain the top $n_{\text{keep}}$ samples ranked by predicted influence magnitude:
$\mathcal{K}_{\text{keep}}=\text{top-}n_{\text{keep}}\{k:|\hat{\theta}_{k}|\}$ (16)
This fixed-size pruning step preserves the samples that matter most for the control update. The MPPI control update is then computed using only the retained samples:
@@
Adaptive Constraint Penalty for Safety. To adapt the constraint penalty online, we monitor the fraction of sampled trajectories that violate constraints. We define the violation rate as
$r_{\text{viol}}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{1}[C_{\text{viol},k}>0]$ (18)
and maintain an exponential moving average $\hat{r}_{\text{viol},t}=(1-\alpha_{\text{ema}})\hat{r}_{\text{viol},t-1}+\alpha_{\text{ema}}\,r_{\text{viol},t}$ with $\alpha_{\text{ema}}=0.1$. When $\hat{r}_{\text{viol}}$ exceeds the target, many trajectories violate constraints, suggesting the penalty $\rho$ is insufficient. We update
$\rho_{t+1}=\text{clip}\left(\rho_{t}+\eta\cdot(\hat{r}_{\text{viol},t}-r_{\text{target}}),\;\rho_{\min},\;\rho_{\max}\right)$ (19)
where $\eta>0$ is a step size and $r_{\text{target}}\in(0,1)$ is the desired maximum violation rate. Algorithm 2 presents the complete online procedure.
@@
0: Trained predictor $h_{\phi^{*}}$, initial state $x_{0}$, nominal control $\bar{U}$, initial penalty $\rho$, number to keep $n_{\text{keep}}$, target violation rate $r_{\text{target}}$, step size $\eta$, EMA parameter $\alpha_{\text{ema}}$, temperature $\lambda$
[…]
2:for each control iteration do
3:  Sample $K$ perturbations $\{\epsilon_{k}\}$, roll out trajectories from current state
4:  Compute cost components: $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ for all $k$
5:  Compute total costs: $C_{k}\leftarrow C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$
6:  Build features $\mathbf{f}_{k}$ and predict influence: $\hat{\theta}_{k}\leftarrow h_{\phi^{*}}(\mathbf{f}_{k})$
7:  Select top $n_{\text{keep}}$ samples by $|\hat{\theta}_{k}|$: $\mathcal{K}_{\text{keep}}$
8:  Compute violation rate: $r_{\text{viol}}\leftarrow\frac{1}{K}\sum_{k}\mathbf{1}[C_{\text{viol},k}>0]$
9:  Update EMA: $\hat{r}_{\text{viol}}\leftarrow(1-\alpha_{\text{ema}})\hat{r}_{\text{viol}}+\alpha_{\text{ema}}\,r_{\text{viol}}$
10:  Update penalty: $\rho\leftarrow\text{clip}(\rho+\eta(\hat{r}_{\text{viol}}-r_{\text{target}}),\rho_{\min},\rho_{\max})$
11:  Recompute weights over $\mathcal{K}_{\text{keep}}$: $\tilde{w}_{k}\leftarrow\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$
12:  Compute control: $u^{*}\leftarrow\bar{u}+\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,\epsilon_{k}$
@@
Remark 1 (Pruning and Penalty Adaptation)
The pruning step and the penalty adaptation work independently of each other. Pruning keeps the samples with the highest predicted influence on the control update, so the weighted average is computed over fewer but more informative samples. The adaptive penalty tracks how often sampled trajectories violate constraints and adjusts $\rho$ to compensate. It is worth noting that pruning by influence magnitude does not automatically favor safe trajectories — a sample with high influence can still violate constraints (see Section IV).
Remark 2 (Computational Complexity)
The offline phase involves $O(N\cdot M\cdot K)$ subset evaluations and $N$ LASSO solves, but it only needs to be done once. At runtime, the online phase only requires $O(K)$ forward passes through $h_{\phi^{*}}$ per control step, which is small compared to the cost of sampling trajectories and evaluating costs. Pruning also reduces the weighted average computation from $O(K)$ down to $O(|\mathcal{K}_{\text{keep}}|)$.
@@ IV Experiments
We evaluate the proposed framework on two closed-loop tracks: a heart-shaped course and a figure-8 course, with static obstacle avoidance using a kinematic bicycle model. All methods share identical dynamics and track geometry; they differ only in sample count and cost/sampling strategy. All experiments use 8 random seeds and report the mean $\pm$ standard deviation.
IV-A Experimental Setup
Dynamics. We use a kinematic bicycle model with state $x=[p_{x},p_{y},\psi,v]^{\top}$ and control $u=[a,\delta]^{\top}$ (acceleration, steering angle):
$p_{x,t+1}$ $=p_{x,t}+v_{t}\cos(\psi_{t})\,\Delta t$
$p_{y,t+1}$ $=p_{y,t}+v_{t}\sin(\psi_{t})\,\Delta t$
$\psi_{t+1}$ $=\psi_{t}+\frac{v_{t}}{L}\tan(\delta_{t})\,\Delta t$
$v_{t+1}$ $=\text{clip}(v_{t}+a_{t}\,\Delta t,\;v_{\min},\;v_{\max})$ (20)
with wheelbase $L=0.3$ m, $\Delta t=0.1$ s, $v_{\max}=2.5$ m/s, $v_{\min}=0.3$ m/s, and maximum steering angle $|\delta|\leq 0.55$ rad.
Track environments. We consider two closed-loop reference paths:
Heart track. A modified parametric heart curve:
$h_{x}(t)$ $=s\cdot 16\sin^{3}(t),$
$h_{y}(t)$ $=s\cdot(13\cos t-3\cos 2t-0.8\cos 3t-0.2\cos 4t)$ (21)
with scale $s=0.6$, centered at $(5,5)$ m. The track has an arc length of $53.3$ m, a constant width of $0.7$ m, and ten rectangular obstacles ($0.40\times 0.28$ m) placed at evenly spaced arc length fractions with alternating lateral offsets.
Lemniscate track. A lemniscate of Bernoulli:
$\ell_{x}(t)$ $=\frac{a\cos t}{1+\sin^{2}t},\qquad\ell_{y}(t)=\frac{a\sin t\cos t}{1+\sin^{2}t}$ (22)
with $a=4.5$, centered at $(5,5)$ m. The figure-8 has a self-intersection where the track narrows to $30\%$ of the base width ($0.7$ m). Eight obstacles are placed along the path with tighter lateral offsets near the crossing.
IV-B Baselines
We compare nine methods spanning standard MPPI at multiple sample budgets, heuristic pruning, and the proposed datamodel-based variants (Fig. 2). MPPI K=50/100/200/500K\!=\!50/100/200/500 runs standard MPPI with i.i.d. samples; K=50K\!=\!50 uses only quadratic centerline cost ($w_{\text{track}}=8.0$), obstacle barriers, and a progress reward, while $K\geq 100$ adds boundary violation and heading alignment penalties. All four pruning baselines start with K=500K\!=\!500 samples and keep 100. Random Pruning simply picks 100 at random, giving a lower bound on what pruning can do. Top-Weight Pruning keeps the 100 samples that already carry the most MPPI importance weight $w_{k}$—in other words, the trajectories that would dominate the weighted average anyway. Cost-Threshold Pruning throws out the 400 most expensive trajectories and recomputes MPPI weights over the remaining 100. Elite-CEM also keeps the 100 cheapest samples, but instead of applying MPPI weights, it just averages their perturbations uniformly, following the Cross-Entropy Method idea of working with an elite set. DM-fixed builds on MPPI K=500K\!=\!500 with temporally correlated AR(1) noise ($\beta=0.6$), 3 MPPI iterations per timestep, control smoothness penalty, curvature-aware target speed, and influence-based pruning. DM-adapted extends DM-fixed with adaptive temperature scaling, log-barrier boundary cost, look-ahead curvature-aware speed, ESS-driven noise adaptation, and elite-sample refinement.
Figure 2: Representative trajectories on the heart track (left) and lemniscate track (right).
IV-C Results
TABLE I: Results across both tracks (8 seeds). Pruning methods draw K=500K\!=\!500 and retain 100 samples.
Method Tracking RMSE Heading RMSE Boundary
(m) (rad) Viol.
Heart track (10 obstacles, arc length 53.3 m)
MPPI K=50K\!=\!50 $0.65\pm 0.10$ $0.39\pm 0.12$ $88.7\pm 12.1$
MPPI K=100K\!=\!100 $8\pm 0.06 0.33±0$$3\pm 0.08$ $31.3\pm 8.5$
MPPI K=200K\!=\!200 $0.25\pm 0.04$ $0.30\pm 0.07$ $6.0\pm 5.2$
MPPI K=500K\!=\!500 $0.18\pm 0.02$ $0.28\pm 0.01$ $0.7\pm 0.6$
Random Prune $0.71\pm 0.06$ $0.42\pm 0.08$ $97.3\pm 14.8$
Top-Weight $0.21\pm 0.03$ $0.26\pm 0.10$ $2.0\pm 1.7$
Cost-Threshold $0.24\pm 0.04$ $0.26\pm 0.08$ $6.0\pm 2.7$
Elite-CEM $0.28\pm 0.06$ $0.26\pm 0.05$ $8.7\pm 1.5$
DM-fixed $0.17\pm 0.02$ $\mathbf{0.23\pm 0.03}$ $1.2\pm 0.4$}
DM-adapted $\mathbf{0.17\pm 0.01}$ $0.24\pm 0.05$ $\mathbf{0.8\pm 0.3}$
Lemniscate track (8 obstacles, self-intersection)
MPPI K=50K\!=\!50 $0.58\pm 0.10$ $0.36\pm 0.03$ $62.3\pm 10.4$
MPPI K=100K\!=\!100 $0.34\pm 0.03$ $0.30\pm 0.02$ $22.7\pm 6.8$
MPPI K=200K\!=\!200 $0.22\pm 0.02$ $0.27\pm 0.02$ $5.3\pm 4.1$
MPPI K=500K\!=\!500 $0.16\pm 0.01$ $0.25\pm 0.01$ $1.3\pm 0.5$
Random Prune $0.64\pm 0.07$ $0.39\pm 0.14$ $71.0\pm 11.6$
Top-Weight $0.25\pm 0.07$ $0.23\pm 0.07$ $8.3\pm 1.2$
Cost-Threshold $0.31\pm 0.05$ $0.23\pm 0.12$ $9.4\pm 4.2$
Elite-CEM $0.39\pm 0.08$ $0.24\pm 0.05$ $7.0\pm 3.0$
DM-fixed $0.19\pm 0.03$ $\mathbf{0.21\pm 0.08}$ $2.1\pm 0.7$}
DM-adapted $\mathbf{0.16\pm 0.02}$ $0.22\pm 0.05$ $\mathbf{1.0\pm 0.5}$
Table I presents results across both tracks. The clearest takeaway is that how you prune matters far more than whether you prune at all. Random pruning actually performs worse than every standard MPPI configuration, including the smallest sample budget K=50K\!=\!50—throwing out samples indiscriminately hurts the weighted average by dropping useful trajectories while keeping noisy ones.
Informed pruning, on the other hand, brings clear improvements. Top-Weight, Cost-Threshold, and Elite-CEM all reduce tracking error compared to unpruned MPPI K=500K\!=\!500, though the gains vary across methods and tracks. All three achieve similar heading RMSE, which is expected since they are all selecting from the same pool of low-cost or high-weight samples. However, these heuristic methods still incur several boundary violations on both tracks, with the numbers getting worse on the lemniscate, where the narrowed self-intersection leaves less margin for error.
DM-fixed and DM-adapted consistently achieve the fewest boundary violations across both tracks. DM-adapted reaches the best or near-best tracking RMSE on both courses while keeping violations roughly an order of magnitude lower than the heuristic pruning baselines. The correlated noise, multi-iteration refinement, and influence-based pruning in DM-fixed produce smoother trajectories that naturally stay further from boundaries, while the adaptive mechanisms in DM-adapted further sharpen tracking by slowing down early in tight sections.
These trends hold across track geometries. The relative ordering of methods stays consistent on both tracks despite their different difficulties: the lemniscate’s self-intersection and variable width demand tighter control, yet the DM-adapted method keeps violations near one on average, while heuristic pruning methods see their violation counts climb.
Finally, simply adding more samples shows diminishing returns—going from K=200K\!=\!200 to K=500K\!=\!500 barely changes tracking RMSE. Yet pruning K=500K\!=\!500 down to 100 well-chosen samples beats the full unpruned budget. A smaller set of good samples outperforms a larger set of unfiltered ones.
V Conclusion
We presented DM-MPPI, a framework that brings the datamodel idea from machine learning into MPPI control. We tested it on a kinematic bicycle model navigating two closed-loop tracks with obstacles. Informed sample pruning—keeping only the top $20\%$ of trajectories—substantially lowers tracking error compared to unpruned MPPI. Random pruning, by contrast, makes things worse than even the smallest sample budget. Top-weight, cost-threshold, and CEM-style pruning all improve tracking, but they still produce noticeably more boundary violations than the datamodel-based variants. DM-adapted consistently achieves the best or near-best tracking accuracy across both tracks while keeping boundary violations far below all other methods. This suggests that influence-aware selection offers a meaningful safety advantage over simple heuristic pruning.
Future work includes: extending evaluation to higher-dimensional systems and dynamic obstacles to test generality; replacing the linear datamodel with nonlinear or attention-based influence predictors that capture sample interactions; developing online adaptation to mitigate training distribution mismatch; integrating with formal safety mechanisms such as Control Barrier Functions [1] or safe importance sampling [4] for hard constraint guaranties; coupling influence prediction with adaptive resampling, where the pruned budget is reallocated to resample in promising control regions; exploring how DM-MPPI can be combined with other MPPI improvements such as covariance adaptation and warm-starting.
Acknowledgment
Claude was used to assist with the language editing of this manuscript.